find the slope and y-intercept of: y=-3x-5

write the equation of the line through (0,3) with slope -4/7 in slope intercept form.
write the equation of the line through (-2,5) with slope 2 in slope- intercept form.

Answer by lynnlo(4176)

(1)slope=========-3
y intercept======-5
(2)x^2-3x=0
(3)x^2-3x-10=0